OpenStudy (anonymous):

One of the factors of g2 – 32g + 256 is: g + 8 g – 8 g – 16 g + 16

OpenStudy (anonymous):

Oh wait lol nvm. I get the question now. Just what is one of the factors.

OpenStudy (anonymous):

do you mean g^2 -32g + 256 ??

OpenStudy (anonymous):

and yeah sry jk. g^2. Anyway, I know it's g-16.

OpenStudy (anonymous):

-16*-16=256 -16+-16=-32

Parth (parthkohli):

Yay :D

Parth (parthkohli):

Are you able to figure it out now?

OpenStudy (anonymous):

Lol thx :). didn't really what the question was asking for until I posted it on OS. Yes, @ParthKohli and the factors for this are (g-16)(g-16)=(g-16)^2 Perfect square trinomial

Parth (parthkohli):

Yep :)

OpenStudy (anonymous):

Check out wolframalpha.com